# Offline-online approximation of multiscale eigenvalue problems with random defects

### v1

@@ 4.2 Main results for the MLOD method
Suppose $\lambda^{(l)}$ and $\tilde{\lambda}^{(l)}_{H,\mathfrak{k}}$ are the $l$-th eigenvalues of (2) and (9) respectively with a multiplicity $r$. If $\norm{\tilde{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{L^{2}(\Omega)}=1$ for each $j=0,\cdots,r-1$, then,
|$\begin{split}\absolutevalue{\frac{\tilde{\lambda}^{(l)}_{H,\mathfrak{k}}-\lambda^{(l)}}{\lambda^{(l)}}}\lesssim g(\mathfrak{k})+H^{2}.\end{split}$
@@
Note that $\frac{\widetilde{a}_{\mathfrak{k}}(\tilde{u}^{(l)}_{H,\mathfrak{k}},\tilde{u}^{(l)}_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(\tilde{u}^{(l)}_{H,\mathfrak{k}},\tilde{u}^{(l)}_{H,\mathfrak{k}})}=\tilde{\lambda}^{(l)}_{H,\mathfrak{k}}$. Also remind that $\lambda^{(l)}$ of (2) is positive. We use Lemma 4.4 with $v_{H,\mathfrak{k}}=\tilde{u}_{H,\mathfrak{k}}^{(l+j)}$ and Lemma 4.1 to obtain that
$\begin{split}\tilde{\lambda}^{(l)}_{H,\mathfrak{k}}-\lambda^{(l)}&=\frac{\widetilde{a}_{\mathfrak{k}}(\tilde{u}^{(l+j)}_{H,\mathfrak{k}},\tilde{u}^{(l+j)}_{H,\mathfrak{k}})}{\widetilde{m}_{\mathfrak{k}}(\tilde{u}^{(l+j)}_{H,\mathfrak{k}},\tilde{u}^{(l+j)}_{H,\mathfrak{k}})}-\lambda^{(l)}\\ &\leq\norm{u^{(l+j)}-\tilde{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}-\lambda^{l}\norm{u^{(l+j)}-\tilde{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}+\norm{(\mathscr{C}-\mathscr{C}_{\mathfrak{k}})\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\hskip 113.81102pt+\tilde{\lambda}^{l}_{H,\mathfrak{k}}\Big{(}{m}_{\mathfrak{k}}(\tilde{u}^{(l+j)}_{H,\mathfrak{k}},\tilde{u}^{(l+j)}_{H,\mathfrak{k}})-\widetilde{m}_{\mathfrak{k}}(\tilde{u}^{(l+j)}_{H,\mathfrak{k}},\tilde{u}^{(l+j)}_{H,\mathfrak{k}})\Big{)}.\end{split}$
Since |$\absolutevalue{\tilde{\lambda}^{(l)}_{H,\mathfrak{k}}}$ | can be bounded above by the absolute continuous eigenvalue, by Theorem 4.5, Theorem 2.2 and Lemma 4.1, we derive that
|$\begin{split}\absolutevalue{\frac{\tilde{\lambda}^{(l)}_{H,\mathfrak{k}}-\lambda^{(l)}}{\lambda^{(l)}}}&\lesssim\bigg{(}H^{2}+g(\mathfrak{k})\bigg{)}^{2}+g(\mathfrak{k})\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}^{2}+H^{2}.\end{split}$ (27)
The assertion follows since $\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}^{2}$ can be bounded above as previously shown. ∎
@@
Given that $g(\mathfrak{k})\approx H^{2}$, if we compare equation (27) with the eigenvalue results form (7), we can clearly see that the reduction of order from $H^{4}$ to $H^{2}$ follows as a consequence of the localization error regardless of the choice of the mass matrix as long as the latter maintains a convergence rate $H^{2}$. Therefore, by choosing the FEM mass matrix instead of a LOD based mass matrix, we obtain an efficient solver at the same convergence rate $H^{2}$.
@@ 4.3 OLOD method
Suppose the OLOD problem in (12) is well-posed. Let $\lambda^{(l)}$ be an eigenvalue of (2) of multiplicity $r$ with the corresponding eigenspace $E_{l+j},\;{j=0,1,\cdots,r-1}$ spanned by the $r$-many orthonormal basis functions. Further let $(\lambda^{(l)},u^{(l+j)})$ and $(\hat{\lambda}^{(l)}_{H,\mathfrak{k}},\hat{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}})$ be the $l$-th solution to (2) and the $l$-th approximated solution to (12) respectively where $\hat{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}=\hat{u}^{(l+j)}_{H,\mathfrak{k}}-\widehat{\mathscr{C}}_{\mathfrak{k}}\hat{u}^{(l+j)}_{H}$, $\hat{u}^{(l+j)}_{H,\mathfrak{k}}\in V_{H}$. For anj $j$, assume the normalization condition $\norm{\hat{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{L^{2}(\Omega)}=1$. Denote $\widehat{E}^{\mathfrak{k}}_{l+j}$ to be the eigenspace spanned by the $r$-many approximated basis functions of (12). If $l+r-1\leq N$, there exists a function ${u}^{(l+j),\mathrm{ms}}_{H}\in{E}_{l+j}$ such that for all $j$ the following holds:
$\norm{u^{(l+j)}-\hat{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}_{\mathcal{A}}\lesssim H^{2}+g(\mathfrak{k})+\Big{(}\max_{T\in\mathcal{T}_{H}}E_{T}\Big{)}\mathfrak{k}^{d/2}.$
@@
$\begin{split}\norm{u^{(l+j)}-\hat{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}_{\mathcal{A}}&\leq\norm{u^{(l+j)}-\tilde{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}_{\mathcal{A}}+\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}-{\mathscr{C}_{\mathfrak{k}}}\tilde{u}^{(l+j)}_{H\mathfrak{k}}-\hat{u}^{(l+j)}_{H,\mathfrak{k}}+\widehat{\mathscr{C}}_{\mathfrak{k}}\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\hskip 86.78099pt+\norm{\widehat{\mathscr{C}}_{\mathfrak{k}}\hat{u}^{(l+j)}_{H,\mathfrak{k}}+{\mathscr{C}_{\mathfrak{k}}}\hat{u}^{(l+j)}_{H,\mathfrak{k}}-{\mathscr{C}_{\mathfrak{k}}}\hat{u}^{(l+j)}_{H,\mathfrak{k}}-{\mathscr{C}}_{\mathfrak{k}}\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\leq\norm{u^{(l+j)}-\tilde{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}_{\mathcal{A}}+\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}-\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\hskip 86.78099pt+\norm{{\mathscr{C}_{\mathfrak{k}}}(\tilde{u}^{(l+j)}_{H,\mathfrak{k}}-\hat{u}^{(l+j)}_{H,\mathfrak{k}})}_{\mathcal{A}}+\norm{(\widehat{\mathscr{C}}_{\mathfrak{k}}-{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\lesssim\norm{u^{(l+j)}-\tilde{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}_{\mathcal{A}}+2\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}-\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}+\norm{(\widehat{\mathscr{C}}_{\mathfrak{k}}-{\mathscr{C}}_{\mathfrak{k}})\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}.\end{split}$
@@
$\begin{split}\norm{\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)}-\widehat{T}^{\mathfrak{k}}_{H}f^{(l+j)}}_{\mathcal{A}}\lesssim&\;\Big{(}\max\limits_{T\in\mathcal{T}_{H}}E_{T}\Big{)}\mathfrak{k}^{d/2}\norm{\widehat{T}^{\mathfrak{k}}_{H}f^{(l+j)}}_{\mathcal{A}}.\end{split}$
Due to the coercivity condition in (15) together with (26) and the interpolation property, we can derive a constant upper bound for $\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}$. Hence, similar arguments used for (25) yield
$\begin{split}\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}-\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}&\lesssim\Big{(}\max\limits_{T\in\mathcal{T}_{H}}E_{T}\Big{)}\mathfrak{k}^{d/2}.\end{split}$ (28)
@@
$\begin{split}\norm{u^{(l+j)}-\hat{u}^{(l+j),ms}_{H,\mathfrak{k}}}_{\mathcal{A}}\lesssim H^{2}+g(\mathfrak{k})+2\Big{(}\max\limits_{T\in\mathcal{T}_{H}}E_{T}\Big{)}\mathfrak{k}^{d/2}+\Big{(}\max_{T\in\mathcal{T}_{H}}E_{T}\Big{)}\mathfrak{k}^{d/2}.\qed\end{split}$
@@
Suppose $\lambda^{(l)}$ and $\hat{\lambda}^{(l)}_{H,\mathfrak{k}}$ are the $l$-th eigenvalues of (2) and (12) respectively with a multiplicity $r$. Consider the same setting as in Theorem 4.8 where $\norm{\hat{u}_{H,\mathfrak{k}}^{(l+j),ms}}_{L^{2}(\Omega)}=1$ for each $j=0,\cdots,r-1$. Then,
|$\begin{split}\absolutevalue{\frac{\hat{\lambda}^{(l)}_{H,\mathfrak{k}}-\lambda^{(l)}}{\lambda^{(l)}}}\lesssim\;&H^{2}+g(\mathfrak{k})+\Big{(}\max\limits_{T\in\mathcal{T}_{H}}E_{T}\Big{)}\mathfrak{k}^{d/2}+\Big{(}\max_{T\in\mathcal{T}_{H}}E_{T}\Big{)}^{2}\mathfrak{k}^{d}.\end{split}$
@@
Note that $\frac{\widehat{a}_{\mathfrak{k}}(\hat{u}^{(l)}_{H,\mathfrak{k}},\hat{u}^{(l)}_{H,\mathfrak{k}})}{\widehat{m}_{\mathfrak{k}}(\hat{u}^{(l)}_{H,\mathfrak{k}},\hat{u}^{(l)}_{H,\mathfrak{k}})}=\hat{\lambda}^{(l)}_{H,\mathfrak{k}}$. Then, we obtain that
$\hat{\lambda}^{(l)}_{H,\mathfrak{k}}-\lambda^{(l)}$ $=\frac{\widehat{a}_{\mathfrak{k}}(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})}{\widehat{m}_{\mathfrak{k}}(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})}-\lambda^{(l)}$
$\leq\widehat{a}_{\mathfrak{k}}(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})-\lambda^{(l)}+\frac{\widehat{a}_{\mathfrak{k}}(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})}{\widehat{m}_{\mathfrak{k}}(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})}-\widehat{a}_{\mathfrak{k}}(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})$
@@
$=(\widehat{a}_{\mathfrak{k}}-\widetilde{a}_{\mathfrak{k}})(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})+[\widetilde{a}_{\mathfrak{k}}(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})-\lambda^{(l)}]+\hat{\lambda}^{(l)}_{H,\mathfrak{k}}\Big{(}\norm{\hat{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}-\widehat{m}_{\mathfrak{k}}(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})\Big{)}$
@@
|$\leq\absolutevalue{(\widehat{a}_{\mathfrak{k}}-\widetilde{a}_{\mathfrak{k}})(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})}+\absolutevalue{\widetilde{a}_{\mathfrak{k}}(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})-\lambda^{(l)}}+\absolutevalue{\widehat{C}_{{\lambda}}\lambda^{(l)}}\,\absolutevalue{\norm{\hat{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}-\widehat{m}_{\mathfrak{k}}(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}})}.$ | . (29)

### v2

@@ 4.2 Main results for the MLOD method
Let $\lambda^{(l)}$ be the $l$-th eigenvalue of the continuous problem (3) with multiplicity $r$, i.e., $\lambda^{(l)}=\lambda^{(l+1)}=\cdots=\lambda^{(l+r-1)}$. Let $\tilde{\lambda}^{(l+j)}_{H,\mathfrak{k}}$ denote the $(l+j)$-th eigenvalue of the discrete multiscale problem (8) for $j=0,1\cdots,r-1$, and assume that $\norm{\tilde{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}}_{L^{2}(\Omega)}=1$ for every $j$. Then, for each $j=0,\cdots,r-1$
|$\begin{split}\absolutevalue{\frac{\tilde{\lambda}^{(l+j)}_{H,\mathfrak{k}}-\lambda^{(l+j)}}{\lambda^{(l+j)}}}\lesssim g(\mathfrak{k})+H^{2}.\end{split}$
@@
Note that $\frac{\widetilde{a}_{\mathfrak{k}}\left(\tilde{u}^{(l+j)}_{H,\mathfrak{k}},\tilde{u}^{(l+j)}_{H,\mathfrak{k}}\right)}{\widetilde{m}_{\mathfrak{k}}\left(\tilde{u}^{(l+j)}_{H,\mathfrak{k}},\tilde{u}^{(l+j)}_{H,\mathfrak{k}}\right)}=\tilde{\lambda}^{(l+j)}_{H,\mathfrak{k}}$. Also remind that $\lambda^{(l+j)}$ of (3) is positive. We use Lemma 4.4 with $v_{H,\mathfrak{k}}=\tilde{u}_{H,\mathfrak{k}}^{(l+j)}$ to obtain that
$\begin{split}\tilde{\lambda}^{(l+j)}_{H,\mathfrak{k}}-\lambda^{(l+j)}&=\frac{\widetilde{a}_{\mathfrak{k}}\left(\tilde{u}^{(l+j)}_{H,\mathfrak{k}},\tilde{u}^{(l+j)}_{H,\mathfrak{k}}\right)}{\widetilde{m}_{\mathfrak{k}}\left(\tilde{u}^{(l+j)}_{H,\mathfrak{k}},\tilde{u}^{(l+j)}_{H,\mathfrak{k}}\right)}-\lambda^{(l+j)}\\ &\leq\norm{u^{(l+j)}-\tilde{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{\mathcal{A}}-\lambda^{(l+j)}\norm{u^{(l+j)}-\tilde{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}+\norm{\left(\mathscr{C}-\mathscr{C}_{\mathfrak{k}}\right)\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\hskip 113.81102pt+\tilde{\lambda}^{(l+j)}_{H,\mathfrak{k}}\left({m}_{\mathfrak{k}}\left(\tilde{u}^{(l+j)}_{H,\mathfrak{k}},\tilde{u}^{(l+j)}_{H,\mathfrak{k}}\right)-\widetilde{m}_{\mathfrak{k}}\left(\tilde{u}^{(l+j)}_{H,\mathfrak{k}},\tilde{u}^{(l+j)}_{H,\mathfrak{k}}\right)\right).\end{split}$
Since |$\absolutevalue{\tilde{\lambda}^{(l+j)}_{H,\mathfrak{k}}}$ | can be bounded above by the absolute continuous eigenvalue, by Theorem 4.5, Theorem 2.2 and Lemma 4.1, we derive that
|$\begin{split}\absolutevalue{\frac{\tilde{\lambda}^{(l+j)}_{H,\mathfrak{k}}-\lambda^{(l+j)}}{\lambda^{(l+j)}}}&\lesssim\left(H^{2}+g(\mathfrak{k})\right)^{2}+g(\mathfrak{k})\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}^{2}+H^{2}.\end{split}$ (24)
The assertion follows since $\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}^{2}$ can be bounded above as previously shown in (23). ∎
@@
Given that $g(\mathfrak{k})\approx H^{2}$, if we compare equation (24) with the eigenvalue results form (7), we can clearly notice that the order of the first term in (24) yields the same order as in (7) whereas the reduction of order from $H^{4}$ to $H^{2}$ follows as a consequence of the error due to the localization corrector term (second term in (24)) in the Petrov-Galerkin method regardless of the choice of the mass matrix. Therefore, by choosing the FEM mass matrix instead of a LOD based mass matrix, we obtain an efficient solver at the same convergence rate $H^{2}$.
@@ 4.3 OLOD method
Suppose the OLOD problem in (10) is well-posed. Let $\lambda^{(l)}$ be the $l$-th eigenvalue of the continuous problem (3) with multiplicity $r$, i.e., $\lambda^{(l)}=\lambda^{(l+1)}=\cdots=\lambda^{(l+r-1)}$, and let the corresponding eigenspace $E^{(l+j)}\subset\mathcal{H}^{1}_{\#}(\Omega)$ for ${j=0,1,\cdots,r-1}$ be spanned by the $r$-many orthonormal basis functions. Let $\left\{\hat{u}_{H,\mathfrak{k}}^{(l+j),\mathrm{ms}}\right\}_{j=0}^{r-1}\subset V^{\mathrm{ms}}_{H,\mathfrak{k}}$ be an orthonormal set of multiscale eigenfunctions approximating the OLOD problem in (10), where $\hat{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}=\hat{u}^{(l+j)}_{H,\mathfrak{k}}-\widehat{\mathscr{C}}_{\mathfrak{k}}\hat{u}^{(l+j)}_{H}$, $\hat{u}^{(l+j)}_{H,\mathfrak{k}}\in V_{H}$. Assume $l+r-1\leq N$, where $N=\mathrm{dim}(V_{H})$. Then, for each $j=0,1,\cdots,r-1$, there exists a function $\hat{u}^{(l+j)}\in E^{(l+j)}$ with $\norm{\hat{u}^{(l+j)}}_{L^{2}(\Omega)}=1$ such that
$\norm{\hat{u}^{(l+j)}-\hat{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}_{\mathcal{A}}\lesssim H^{2}+g(\mathfrak{k})+\left(\max_{T\in\mathcal{T}_{H}}E_{T}\right)\mathfrak{k}^{d/2}.$
The hidden constant depends only on the spectral bounds, domain $\Omega$, multiplicity $r$ and the corresponding continuous eigenvalue $\lambda^{(l)}$.
In the following proof, we use similar arguments used in Theorem 4.5 for the validity of the traingle inequality and the estimates used for bounding $\norm{\hat{u}^{(l+j)}-\tilde{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}_{\mathcal{A}}$.
@@
$\begin{split}\norm{\hat{u}^{(l+j)}-\hat{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}_{\mathcal{A}}&\leq\norm{\hat{u}^{(l+j)}-\tilde{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}_{\mathcal{A}}+\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}-{\mathscr{C}_{\mathfrak{k}}}\tilde{u}^{(l+j)}_{H\mathfrak{k}}-\hat{u}^{(l+j)}_{H,\mathfrak{k}}+\widehat{\mathscr{C}}_{\mathfrak{k}}\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\leq\norm{\hat{u}^{(l+j)}-\tilde{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}_{\mathcal{A}}+\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}-\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\hskip 86.78099pt+\norm{\widehat{\mathscr{C}}_{\mathfrak{k}}\hat{u}^{(l+j)}_{H,\mathfrak{k}}+{\mathscr{C}_{\mathfrak{k}}}\hat{u}^{(l+j)}_{H,\mathfrak{k}}-{\mathscr{C}_{\mathfrak{k}}}\hat{u}^{(l+j)}_{H,\mathfrak{k}}-{\mathscr{C}}_{\mathfrak{k}}\tilde{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\leq\norm{\hat{u}^{(l+j)}-\tilde{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}_{\mathcal{A}}+\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}-\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\hskip 86.78099pt+\norm{{\mathscr{C}_{\mathfrak{k}}}\left(\tilde{u}^{(l+j)}_{H,\mathfrak{k}}-\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)}_{\mathcal{A}}+\norm{\left(\widehat{\mathscr{C}}_{\mathfrak{k}}-{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}\\ &\lesssim\norm{\hat{u}^{(l+j)}-\tilde{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}_{\mathcal{A}}+2\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}-\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}+\norm{\left(\widehat{\mathscr{C}}_{\mathfrak{k}}-{\mathscr{C}}_{\mathfrak{k}}\right)\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}.\end{split}$
@@
$\begin{split}\norm{\widetilde{T}^{\mathfrak{k}}_{H}f^{(l+j)}-\widehat{T}^{\mathfrak{k}}_{H}f^{(l+j)}}_{\mathcal{A}}\lesssim&\;\left(\max\limits_{T\in\mathcal{T}_{H}}E_{T}\right)\mathfrak{k}^{d/2}\norm{\widehat{T}^{\mathfrak{k}}_{H}f^{(l+j)}}_{\mathcal{A}}.\end{split}$
From the coercivity condition in (13) and bounds similar to (23), together with the interpolation property we derive a constant upper bound for $\norm{\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}$. Hence, by similar arguments from compact operator theory used to obtain (22) we yield
$\begin{split}\norm{\tilde{u}^{(l+j)}_{H,\mathfrak{k}}-\hat{u}^{(l+j)}_{H,\mathfrak{k}}}_{\mathcal{A}}&\lesssim\left(\max\limits_{T\in\mathcal{T}_{H}}E_{T}\right)\mathfrak{k}^{d/2}.\end{split}$ (25)
@@
$\begin{split}\norm{\hat{u}^{(l+j)}-\hat{u}^{(l+j),\rm{ms}}_{H,\mathfrak{k}}}_{\mathcal{A}}\lesssim H^{2}+g(\mathfrak{k})+2\left(\max\limits_{T\in\mathcal{T}_{H}}E_{T}\right)\mathfrak{k}^{d/2}+\left(\max_{T\in\mathcal{T}_{H}}E_{T}\right)\mathfrak{k}^{d/2}.\qed\end{split}$
@@
Let $\lambda^{(l)}$ be the $l$-th eigenvalue of the continuous problem (3) with multiplicity $r$, i.e., $\lambda^{(l)}=\lambda^{(l+1)}=\cdots=\lambda^{(l+r-1)}$. Let $\hat{\lambda}^{(l+j)}_{H,\mathfrak{k}}$ denote the $(l+j)$-th eigenvalue of the discrete multiscale problem given by (10) for $j=0,\cdots,r-1$. Consider the same setting as in Theorem 4.8 where $\norm{\hat{u}^{(l+j),\rm{ms}}_{H,\mathfrak{k}}}_{L^{2}(\Omega)}=1$ for every $j$. Then, for each $j=0,\cdots,r-1$
|$\begin{split}\absolutevalue{\frac{\hat{\lambda}^{(l+j)}_{H,\mathfrak{k}}-\lambda^{(l+j)}}{\lambda^{(l+j)}}}\lesssim\;&H^{2}+g(\mathfrak{k})+\left(\max\limits_{T\in\mathcal{T}_{H}}E_{T}\right)\mathfrak{k}^{d/2}+\left(\max_{T\in\mathcal{T}_{H}}E_{T}\right)^{2}\mathfrak{k}^{d}.\end{split}$
@@
Note that $\frac{\widehat{a}_{\mathfrak{k}}\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)}{\widehat{m}_{\mathfrak{k}}\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)}=\hat{\lambda}^{(l+j)}_{H,\mathfrak{k}}$. Then, we obtain that
$\hat{\lambda}^{(l+j)}_{H,\mathfrak{k}}-\lambda^{(l+j)}$ $=\frac{\widehat{a}_{\mathfrak{k}}\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)}{\widehat{m}_{\mathfrak{k}}\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)}-\lambda^{(l+j)}$
$\leq\widehat{a}_{\mathfrak{k}}\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)-\lambda^{(l+j)}+\frac{\widehat{a}_{\mathfrak{k}}\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)}{\widehat{m}_{\mathfrak{k}}\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)}-\widehat{a}_{\mathfrak{k}}\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)$
$=\left(\widehat{a}_{\mathfrak{k}}-\widetilde{a}_{\mathfrak{k}}\right)\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)+\left(\widetilde{a}_{\mathfrak{k}}\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)-\lambda^{(l+j)}\right)$
@@
$\hskip 120.07066pt+\hat{\lambda}^{(l+j)}_{H,\mathfrak{k}}\left(\norm{\hat{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}-\widehat{m}_{\mathfrak{k}}\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)\right)$
|$\leq\absolutevalue{\left(\widehat{a}_{\mathfrak{k}}-\widetilde{a}_{\mathfrak{k}}\right)\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)}+\absolutevalue{\widetilde{a}_{\mathfrak{k}}\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)-\lambda^{(l+j)}}$ |
@@
+|$\hskip 126.61467pt+\absolutevalue{\widehat{C}_{{\lambda}}\lambda^{(l+j)}}\,\absolutevalue{\norm{\hat{u}^{(l+j),\mathrm{ms}}_{H,\mathfrak{k}}}^{2}_{L^{2}(\Omega)}-\widehat{m}_{\mathfrak{k}}\left(\hat{u}^{(l+j)}_{H,\mathfrak{k}},\hat{u}^{(l+j)}_{H,\mathfrak{k}}\right)}$ | (26)